OpenStudy (anonymous):

y = (e^x + 10)/(e^x - 10). Find (dy)/(dx) im confused w/ the quotient rule

OpenStudy (anonymous):

heres what i have found

OpenStudy (anonymous):

(e^x+10)1-(e^x+10)(1)/(e^x-10)^2

OpenStudy (anonymous):

is that right so far?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

quotient rule of f(x)/g(x) is [g(x)*f'(x)-f(x)*g'(x)]/g^2(x)

OpenStudy (anonymous):

i know (g(x)f^'(x)-f(x)g^'(x))/([g(x)]^2).

OpenStudy (anonymous):

dy/dx of (e^x+10) is not 1

OpenStudy (anonymous):

e^x =1 right and the deriv of 10 =0

OpenStudy (anonymous):

no e^x = e^x

OpenStudy (anonymous):

i thought e^x was number

OpenStudy (anonymous):

it is but you are differentiating x not e

OpenStudy (anonymous):

so its (e^x+10)(e^x)-(e^x+10)(e^x)/(e^x-10)^2

OpenStudy (anonymous):

the left side the first term is e^x-10

OpenStudy (anonymous):

ok how do i reduce it?

OpenStudy (anonymous):

factoring out what?

OpenStudy (anonymous):

just multiply through and see is anything drops out

OpenStudy (anonymous):

i dont factor out e^x?

OpenStudy (anonymous):

why would you it doesnt help reduce it any

OpenStudy (anonymous):

oh ok i got it

OpenStudy (anonymous):

it is (e^x)(-20)/(e^x-10)^2

OpenStudy (anonymous):

yes it is, great work